Duffy Enterprises has issued common stock, preferred stock, and debt to finance its operation. The table below shows financial information on the firm.

2018 Balance Sheet:
Interest bearing debt $630,655.00
Preferred stock $48,404.00
Shareholder equity $750,000.00

Duffy has publicly traded common stock that has a capitalization currently of $1,759,233.00.

An analyst has estimated that the cost of debt for Duffy is 7.00%, the cost of the preferred stock is 10.00%, and the cost of the common equity is 13.00%. Duffy has a marginal tax rate of 37.00%.

What is the cost of capital for Duffy?

Answer Format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924))
Answer #1

Common Stock Value = E = $ 1759233, Cost of Common Stock = ke = 13 %, Debt Value = Interest Bearing Debt = D = $ 630655, Preferred Stock Value = P = $ 48404, Cost of Preferred Stock = kp = 10 % and Cost of Debt = kd = 7%

Tax Rate = t = 37%

Total Value = E + D + P = 1759233 + 630655 + 48404 = $ 2438292

Cost of Capital = ke x (E/V) + kd x (1-t) x (D/V) + kp x (P/V) = 13 x (1759233/2438292) + 7 x (1-0.37) x (630655/2438292) + 10 x (48404/2438292) = 10.7187 % ~ 10.72 %